### v2

# Tuning LLMs with Contrastive Alignment Instructions for Machine Translation in Unseen, Low-resource Languages

Abstract
This article introduces contrastive alignment instructions (AlignInstruct) to address two challenges in machine translation (MT) on large language models (LLMs). One is the expansion of supported languages to previously unseen ones. The second relates to the lack of data in low-resource languages. Model fine-tuning through MT instructions (MTInstruct) is a straightforward approach to the first challenge. However, MTInstruct is limited by weak cross-lingual signals inherent in the second challenge. AlignInstruct emphasizes cross-lingual supervision via a cross-lingual discriminator built using statistical word alignments. Our results based on fine-tuning the BLOOMZ models (1b1, 3b, and 7b1) in up to 24 unseen languages showed that: (1) LLMs can effectively translate unseen languages using MTInstruct; (2) AlignInstruct led to consistent improvements in translation quality across 48 translation directions involving English; (3) Discriminator-based instructions outperformed their generative counterparts as cross-lingual instructions; (4) AlignInstruct improved performance in 30 zero-shot directions.
Tuning LLMs with Contrastive Alignment Instructions for Machine Translation in Unseen, Low-resource Languages
Zhuoyuan Mao††thanks: Currently at Sony Group Corporation. Work done during Apple internship. and Yen Yu Apple kevinmzy@gmail.com, yen_yu@apple.com
1 Introduction
Large language models (LLMs) Brown et al. (2020); Chowdhery et al. (2022); Scao et al. (2022); Touvron et al. (2023a); Muennighoff et al. (2023); OpenAI (2023); Anil et al. (2023); Touvron et al. (2023b) achieved good performance for a wide range of NLP tasks for prevalent languages. However, insufficient coverage for low-resource languages remains to be one significant limitation. Low-resource languages are either not present, or orders of magnitude smaller in size than dominant languages in the pre-training dataset. This limitation is in part due to the prohibitive cost incurred by curating good quality and adequately sized datasets for pre-training. Incrementally adapting existing multilingual LLMs to incorporate an unseen, low-resource language thus becomes a cost-effective priority to address this limitation. Previous study de la Rosa and Fernández (2022); Müller and Laurent (2022); Yong et al. (2023) explored extending language support using either continual pre-training Neubig and Hu (2018); Artetxe et al. (2020); Muller et al. (2021); Ebrahimi and Kann (2021), or parameter efficient fine-tuning (PEFT) methods Pfeiffer et al. (2020); Hu et al. (2022); Liu et al. (2022) on monolingual tasks. Extending language support for cross-lingual tasks remains underexplored due to the challenge of incrementally inducing cross-lingual understanding and generation abilities in LLMs Yong et al. (2023).
This study focused on machine translation (MT) to highlight the cross-lingual LLM adaptation challenge. The challenge lies in enabling translation for low-resource languages that often lack robust cross-lingual signals. We first explored the efficacy of fine-tuning LLMs with MT instructions (MTInstruct) in unseen, low-resource languages. MTInstruct is a method previously shown to bolster the translation proficiency of LLMs for supported languages Li et al. (2023). Subsequently, given that cross-lingual alignments are suboptimal in LLMs as a result of data scarcity of low-resource languages, we proposed contrastive alignment instructions (AlignInstruct) to explicitly provide cross-lingual supervision during MT fine-tuning. AlignInstruct is a cross-lingual discriminator formulated using statistical word alignments. Our approach was inspired by prior studies Lambert et al. (2012); Ren et al. (2019); Lin et al. (2020); Mao et al. (2022), which indicated the utility of word alignments in enhancing MT. In addition to AlignInstruct, we discussed two word-level cross-lingual instruction alternatives cast as generative tasks for comparison with AlignInstruct.
Our experiments fine-tuned the BLOOMZ models Muennighoff et al. (2023) of varying sizes (1b1, 3b, and 7b1) for 24 unseen, low-resource languages, and evaluated translation on OPUS-100 Zhang et al. (2020) and Flores-200 Costa-jussà et al. (2022). We first showed that MTInstruct effectively induced the translation capabilities of LLMs for these languages. Building on the MTInstruct baseline, the multi-task learning combining AlignInstruct and MTInstruct resulted in stronger translation performance without the need for additional training corpora. The performance improved with larger BLOOMZ models, as illustrated in Fig. 1, indicating that AlignInstruct is particularly beneficial for larger LLMs during MT fine-tuning. When compared with the generative variants of AlignInstruct, our results indicated that discriminative instructions better complemented MTInstruct. Furthermore, merging AlignInstruct with its generative counterparts did not further improve translation quality, underscoring the efficacy and sufficiency of AlignInstruct in leveraging word alignments for MT.
In zero-shot translation evaluation on the OPUS benchmark, AlignInstruct exhibited improvements over the MTInstruct baseline in 30 zero-shot directions between non-English languages, when exclusively fine-tuned with three unseen languages (German, Dutch, and Russian). However, when incorporating supported languages (Arabic, French, and Chinese) the benefits of AlignInstruct were only evident in zero-shot translations where the target language was a supported language. In addition, to interpret the inherent modifications within the BLOOMZ models after applying MTInstruct or AlignInstruct, we conducted a visualization of the layer-wise cross-lingual alignment capabilities of the model representations.
2 Methodology
This section presents MTInstruct as the baseline, and AlignInstruct. The MTInstruct baseline involved fine-tuning LLMs using MT instructions. AlignInstruct dealt with the lack of cross-lingual signals stemming from the limited parallel training data in low-resource languages. The expectation was enhanced cross-lingual supervision cast as a discriminative task without extra training corpora. Following this, we introduced two generative variants of AlignInstruct for comparison.111We also discussed monolingual instructions for MT fine-tuning in App. F.
2.1 Baseline: MTInstruct
Instruction tuning Wang et al. (2022); Mishra et al. (2022); Chung et al. (2022); Ouyang et al. (2022); Sanh et al. (2022); Wei et al. (2022) has been shown to generalize LLMs’ ability to perform various downstream tasks, including MT Li et al. (2023).
Given a pair of the parallel sentences, , where , . are members of the vocabulary containing unique tokens that accommodate languages and . Li et al. (2023) showed that the following MT instructions (MTInstruct) can improve the translation ability in an LLM with a limited number of parallel sentences:
Input: “Translate from to .
Output: “.”
Note that Li et al. (2023) demonstrated the utility of MTInstruct solely within the context of fine-tuning for languages acquired at pre-training phase. This study called for an assessment of MTInstruct on its efficacy for adapting to previously unsupported languages, denoted as , accompanied by the parallel data in a supported language .
2.2 AlignInstruct
Word alignments have been demonstrated to enhance MT performance Lambert et al. (2012); Ren et al. (2019); Lin et al. (2020); Mao et al. (2022), both in the fields of statistical machine translation (SMT) Brown et al. (1993) and neural machine translation (NMT) Sutskever et al. (2014); Bahdanau et al. (2015). Ren et al. (2019) and Mao et al. (2022) reported the utility of SMT-derived contrastive word alignments in guiding encoder-decoder NMT model training. Built upon their findings, we introduced AlignInstruct for bolstering cross-lingual alignments in LLMs. We expected AlignInstruct to enhancing translation performance particularly for languages with no pre-training data and limited fine-tuning data.
As shown in Fig. 2, we employed FastAlign Dyer et al. (2013) to extract statistical word alignments from parallel corpora. Our approach depended on a trained FastAlign model (IBM Model 2, Brown et al., 1993) to ensure the quality of the extracted word pairs. These high-quality word alignment pairs were regarded as “gold” word pairs for constructing AlignInstruct instructions.222Note that these word pairs may not necessarily represent direct translations of each other; instead, they are word pairs identified based on their co-occurrence probability within the similar context. Refer to IBM model 2 in SMT. Assuming one gold word pair was provided for the sentence pair , the AlignInstruct instruction reads:
Input: “Given the following parallel sentence between and , judge whether the assertion is True or False.
Assertion: “” can be aligned with “” statistically.”
Output: “True” (or “False”)
Instructions with the “False” output were constructed by uniformly swapping out part of the word pair to create misalignment. We anticipated that this treatment forced the model to learn to infer the output by recognizing true alignment-enriched instructions. This would require the model to encode word-level cross-lingual representation, a crucial characteristic for MT tasks.
2.3 Generative Counterparts of AlignInstruct
Previous studies Liang et al. (2022); Yu et al. (2023) have suggested the importance of both discriminative and generative tasks in fine-tuning LLMs. We accordingly considered two generative variants of AlignInstruct. We then compared them with AlignInstruct to determine the most effective training task. As detailed in Sec. 4, our results indicated that these variants underperformed AlignInstruct when applied to unseen, low-resource languages.
2.3.1 HintInstruct
HintInstruct as a generative variant of AlignInstruct was instructions containing word alignment hints. It was inspired by Ghazvininejad et al. (2023), where dictionary hints were shown to improve few-shot in-context leaning. Instead of relying on additional dictionaries, we used the same word alignments described in Sec. 2.2, which were motivated by the common unavailability of high-quality dictionaries for unseen, low-resource languages. Let be word pairs extracted from the sentence pair . HintInstruct follows the instruction pattern:
Input: “Use the following alignment hints and translate from to .
Alignments between and :
Output: “.”
where denotes the number of the word alignment pairs used to compose the instructions. Different from AlignInstruct, HintInstruct expects the translation targets to be generated.
2.3.2 ReviseInstruct
ReviseInstruct was inspired by Ren et al. (2019) and Liu et al. (2020) for the notion of generating parallel words or phrases, thereby encouraging a model to encode cross-lingual alignments. A ReviseInstruct instruction contained a partially corrupted translation target, as well as a directive to identify and revise these erroneous tokens. Tokens are intentionally corrupted at the granularity of individual words, aligning with the word-level granularity in AlignInstruct and HintInstruct. ReviseInstruct follows the instruction pattern:333We illustrated examples of HintInstruct and ReviseInstruct in App. E for reference.
Input: “Given the following translation of from , output the incorrectly translated word and correct it.
Output: “The incorrectly translated word is "". It should be "".”
3 Experimental Settings
| BLOOMZ | Objective | OPUS enxx | OPUS xxen | Flores enxx | Flores xxen | ||||||||
| model | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | |
| BLOOMZ-7b1 | w/o fine-tuning | 3.61 | 8.82 | 47.81 | 6.70 | 18.49 | 51.68 | 2.00 | 9.35 | 36.54 | 9.95 | 24.47 | 52.05 |
| Individual objectives | |||||||||||||
| MTInstruct | 11.54 | 25.33 | 64.54 | 18.59 | 33.25 | 69.84 | 3.30 | 17.10 | 40.58 | 11.37 | 27.14 | 56.33 | |
| AlignInstruct | 4.73 | 9.23 | 49.85 | 5.32 | 12.90 | 53.26 | 1.97 | 8.90 | 42.35 | 3.47 | 11.93 | 39.58 | |
| Multiple objectives with different curricula | |||||||||||||
| MT+Align | 12.28 | 26.17 | 65.54 | 18.72 | 34.02 | 70.69 | 3.26 | 17.20 | 41.07 | 11.60 | 27.38 | 56.98 | |
| AlignMT | 11.73 | 25.48 | 64.54 | 17.54 | 32.62 | 69.76 | 3.35 | 17.21 | 40.85 | 11.32 | 27.21 | 56.50 | |
| MT+AlignMT | 12.10 | 26.16 | 65.43 | 18.23 | 33.54 | 70.60 | 3.28 | 17.26 | 41.13 | 11.48 | 27.34 | 56.78 | |
| BLOOMZ-3b | w/o fine-tuning | 4.63 | 9.93 | 48.53 | 5.90 | 16.38 | 48.05 | 2.00 | 9.09 | 39.52 | 5.86 | 18.56 | 47.03 |
| Individual objectives | |||||||||||||
| MTInstruct | 10.40 | 23.08 | 62.28 | 16.10 | 31.15 | 68.36 | 2.85 | 16.23 | 39.21 | 8.92 | 24.57 | 53.33 | |
| AlignInstruct | 1.70 | 4.05 | 43.89 | 0.87 | 3.20 | 41.93 | 0.16 | 3.09 | 31.10 | 0.10 | 1.80 | 29.46 | |
| Multiple objectives with different curricula | |||||||||||||
| MT+Align | 10.61 | 23.64 | 62.84 | 16.73 | 31.51 | 68.52 | 2.95 | 16.62 | 39.83 | 9.50 | 25.16 | 54.35 | |
| AlignMT | 10.22 | 22.53 | 61.99 | 15.90 | 30.31 | 67.79 | 3.02 | 16.43 | 39.46 | 9.07 | 24.70 | 53.71 | |
| MT+AlignMT | 10.60 | 23.35 | 62.69 | 16.58 | 31.64 | 68.98 | 2.93 | 16.57 | 39.78 | 9.41 | 25.08 | 54.13 | |
| BLOOMZ-1b1 | w/o fine-tuning | 3.76 | 7.57 | 46.98 | 4.78 | 14.11 | 49.34 | 1.24 | 6.93 | 38.13 | 3.49 | 14.56 | 43.26 |
| Individual objectives | |||||||||||||
| MTInstruct | 7.42 | 17.85 | 57.53 | 11.99 | 25.59 | 63.93 | 2.11 | 14.40 | 36.35 | 5.33 | 20.65 | 48.83 | |
| AlignInstruct | 2.51 | 5.29 | 45.17 | 3.13 | 8.92 | 48.48 | 0.35 | 3.79 | 31.70 | 1.35 | 6.43 | 33.63 | |
| Multiple objectives with different curricula | |||||||||||||
| MT+Align | 7.80 | 18.48 | 57.77 | 12.57 | 25.92 | 64.03 | 2.16 | 14.54 | 37.05 | 5.46 | 20.90 | 49.31 | |
| AlignMT | 7.49 | 18.09 | 57.67 | 11.80 | 24.70 | 63.29 | 2.08 | 14.28 | 36.61 | 5.24 | 20.53 | 48.76 | |
| MT+AlignMT | 7.98 | 18.61 | 57.94 | 12.43 | 25.78 | 63.93 | 2.16 | 14.46 | 37.02 | 5.37 | 20.67 | 49.01 | |
3.1 Backbone Models and Unseen Languages
Our experiments fine-tuned the BLOOMZ models Muennighoff et al. (2023) for MT in unseen, low-resource languages. BLOOMZ is an instruction fine-tuned multilingual LLM from BLOOM Scao et al. (2022) that supports translation across languages. Two lines of experiments evaluated the effectiveness of the MTInstruct baseline and AlignInstruct:
BLOOMZ+24 Tuning BLOOMZ-7b1, BLOOMZ-3b, and BLOOMZ-1b1444https://huggingface.co/bigscience/bloomz for unseen, low-resource languages. These experiments aimed to: (1) assess the effectiveness of AlignInstruct in multilingual, low-resource scenarios; (2) offer comparison across various model sizes. We used the OPUS-100 Zhang et al. (2020)555https://opus.nlpl.eu/opus-100.php datasets as training data. OPUS-100 is an English-centric parallel corpora, with around M parallel sentences in total for 24 selected languages, averaging k sentence pairs for each language and English. Refer to App. A for training data statistics. We used OPUS-100 and Flores-200 Costa-jussà et al. (2022)666https://github.com/facebookresearch/flores/blob/main/flores200/README.md for evaluating translation between English and unseen languages (48 directions in total) on in-domain and out-of-domain test sets, respectively. The identical prompt as introduced in Sec. 2.1 was employed for inference. Inferences using alternative MT prompts are discussed in App. G.
BLOOMZ+3 Tuning BLOOMZ-7b1 with three unseen languages, German, Dutch, and Russian, or a combination of these three unseen languages and another three seen (Arabic, French, and Chinese). We denote the respective setting as de-nl-ru and ar-de-fr-nl-ru-zh. These experiments assessed the efficacy of AlignInstruct in zero-shot translation scenarios, where translation directions were not presented during fine-tuning, as well as the translation performance when incorporating supported languages as either source or target languages. To simulate the low-resource fine-tuning scenario, we randomly sampled k parallel sentences for each language. For evaluation, we used the OPUS-100 supervised and zero-shot test sets, comprising 12 supervised directions involving English and 30 zero-shot directions without English among six languages.
Notably, BLOOMZ’s pre-training data includes the English portion of the Flores-200 dataset, potentially leading to data leakage during evaluation Muennighoff et al. (2023); Zhu et al. (2023a). To mitigate this, our evaluation also compared translation quality before and after fine-tuning, thereby distinguishing the genuine improvements in translation capability attributable to the fine-tuning process (refer to the results in Sec. 4).
3.2 Training Details and Curricula
The PEFT method, LoRA Hu et al. (2022), was chosen to satisfy the parameter efficiency requirement for low-resource languages, as full-parameter fine-tuning would likely under-specify the models.See App. B for implementation details. How AlignInstruct and MTInstruct are integrated into training remained undetermined. To that end, we investigated three training curricula:
| Objective | en-af | af-en | en-am | am-en | en-be | be-en | en-cy | cy-en | en-ga | ga-en | en-gd | gd-en |
| MTInstruct | 25.0 | 38.5 | 3.0 | 3.4 | 8.9 | 14.0 | 20.2 | 33.2 | 15.6 | 29.2 | 13.1 | 66.0 |
| MT+Align | 25.0 | 36.9 | 3.4 | 4.9 | 8.3 | 13.9 | 20.6 | 33.8 | 17.6 | 32.6 | 15.6 | 48.1 |
| Objective | en-gl | gl-en | en-ha | ha-en | en-ka | ka-en | en-kk | kk-en | en-km | km-en | en-ky | ky-en |
| MTInstruct | 16.9 | 24.7 | 12.3 | 10.0 | 4.6 | 10.0 | 12.6 | 14.6 | 19.7 | 13.9 | 16.0 | 21.1 |
| MT+Align | 17.1 | 24.4 | 14.6 | 11.4 | 4.9 | 10.5 | 12.3 | 15.6 | 20.4 | 14.4 | 15.8 | 23.3 |
| Objective | en-li | li-en | en-my | my-en | en-nb | nb-en | en-nn | nn-en | en-oc | oc-en | en-si | si-en |
| MTInstruct | 13.5 | 21.3 | 6.2 | 5.2 | 12.7 | 22.2 | 18.3 | 27.1 | 10.0 | 13.4 | 5.2 | 11.5 |
| MT+Align | 13.2 | 22.3 | 7.6 | 6.3 | 13.5 | 24.2 | 19.0 | 28.5 | 9.1 | 13.5 | 5.1 | 13.9 |
| Objective | en-tg | tg-en | en-tk | tk-en | en-tt | tt-en | en-ug | ug-en | en-uz | uz-en | en-yi | yi-en |
| MTInstruct | 5.5 | 8.0 | 24.4 | 30.4 | 1.9 | 3.6 | 1.2 | 4.2 | 3.1 | 5.7 | 7.1 | 14.9 |
| MT+Align | 6.6 | 8.8 | 27.2 | 31.2 | 2.1 | 5.0 | 1.1 | 5.5 | 3.5 | 7.4 | 11.1 | 12.8 |
| BLOOMZ | Objective | OPUS enxx | OPUS xxen | Flores enxx | Flores xxen | ||||||||
| model | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | |
| BLOOMZ-7b1 | MTInstruct | 11.54 | 25.33 | 64.54 | 18.59 | 33.25 | 69.84 | 3.30 | 17.10 | 40.58 | 11.37 | 27.14 | 56.33 |
| MT+Align | 12.28 | 26.17 | 65.54 | 18.72 | 34.02 | 70.69 | 3.26 | 17.20 | 41.07 | 11.60 | 27.38 | 56.98 | |
| MT+Hint | 12.12 | 25.92 | 64.60 | 18.25 | 33.18 | 70.31 | 3.34 | 17.13 | 41.10 | 11.45 | 27.37 | 56.86 | |
| MT+Revise | 11.96 | 25.73 | 64.73 | 18.69 | 33.74 | 70.32 | 3.34 | 17.10 | 41.07 | 11.44 | 27.37 | 56.73 | |
| BLOOMZ-3b | MTInstruct | 10.40 | 23.08 | 62.28 | 16.10 | 31.15 | 68.36 | 2.85 | 16.23 | 39.21 | 8.92 | 24.57 | 53.33 |
| MT+Align | 10.61 | 23.64 | 62.84 | 16.73 | 31.51 | 68.52 | 2.95 | 16.62 | 39.83 | 9.50 | 25.16 | 54.35 | |
| MT+Hint | 10.49 | 23.34 | 62.65 | 16.29 | 31.43 | 68.83 | 3.11 | 16.95 | 39.91 | 9.52 | 25.25 | 54.28 | |
| MT+Revise | 10.52 | 23.03 | 62.04 | 16.22 | 30.98 | 68.28 | 2.99 | 16.83 | 39.52 | 9.47 | 25.21 | 53.91 | |
| BLOOMZ-1b1 | MTInstruct | 7.42 | 17.85 | 57.53 | 11.99 | 25.59 | 63.93 | 2.11 | 14.40 | 36.35 | 5.33 | 20.65 | 48.83 |
| MT+Align | 7.80 | 18.48 | 57.77 | 12.57 | 25.92 | 64.03 | 2.16 | 14.54 | 37.05 | 5.46 | 20.90 | 49.31 | |
| MT+Hint | 7.71 | 18.15 | 57.76 | 11.52 | 24.88 | 63.63 | 2.21 | 14.61 | 37.24 | 5.47 | 20.78 | 48.97 | |
| MT+Revise | 7.31 | 17.99 | 57.45 | 12.00 | 25.33 | 63.81 | 2.07 | 14.32 | 36.68 | 5.41 | 20.91 | 49.09 | |
| Objective | OPUS enxx | OPUS xxen | Flores enxx | Flores xxen | ||||||||
| BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | |
| MTInstruct | 11.54 | 25.33 | 64.54 | 18.59 | 33.25 | 69.84 | 3.30 | 17.10 | 40.58 | 11.37 | 27.14 | 56.33 |
| MT+Align | 12.28 | 26.17 | 65.54 | 18.72 | 34.02 | 70.69 | 3.26 | 17.20 | 41.07 | 11.60 | 27.38 | 56.98 |
| MT+Align+Revise | 12.08 | 25.73 | 64.55 | 19.23 | 34.32 | 70.60 | 3.33 | 17.25 | 41.17 | 11.60 | 27.61 | 57.22 |
| MT+Align+Hint | 12.02 | 25.51 | 64.58 | 19.40 | 34.44 | 70.65 | 3.25 | 16.87 | 41.13 | 11.58 | 27.48 | 56.93 |
| MT+Hint+Revise | 12.10 | 25.69 | 64.68 | 19.58 | 34.49 | 70.55 | 3.34 | 17.24 | 41.13 | 11.70 | 27.62 | 57.19 |
| MT+Align+Hint+Revise | 12.00 | 25.39 | 64.55 | 19.68 | 34.48 | 70.64 | 3.40 | 17.17 | 41.21 | 11.67 | 27.54 | 57.16 |
Multi-task Fine-tuning combined multiple tasks in a single training session Caruana (1997). This was realized by joining MTInstruct and AlignInstruct training data, denoted as MT+Align.777Note that AlignInstruct and MTInstruct were derived from the same parallel corpora.
Pre-fine-tuning & Fine-tuning arranges fine-tuning in a two-stage curriculum Bengio et al. (2009), first with AlignInstruct, then with MTInstruct.888An effective curriculum often starts with a simple and general task, followed by a task-specific task. This configuration, denoted as AlignMT, validates whether AlignInstruct should precede MTInstruct.
Mixed Fine-tuning Chu et al. (2017) arranged the above curricula to start with MT+Align, followed by MTInstruct, denoted as MT+AlignMT.
| Fine-tuned | Objective | Zero-shot Directions | Supervised Directions | ||||||
| Languages | Directions | BLEU | chrF++ | COMET | Directions | BLEU | chrF++ | COMET | |
| - | w/o fine-tuning | overall | 6.89 | 19.14 | 57.95 | enxx | 13.38 | 26.65 | 64.28 |
| xxen | 21.70 | 42.05 | 72.72 | ||||||
| seenseen | 16.95 | 30.78 | 74.58 | enseen | 20.13 | 32.87 | 76.99 | ||
| seenunseen | 2.30 | 13.31 | 49.98 | enunseen | 6.63 | 20.43 | 51.56 | ||
| unseenseen | 7.78 | 20.07 | 62.74 | seenen | 26.30 | 48.70 | 78.22 | ||
| unseenunseen | 2.37 | 14.83 | 46.06 | unseenen | 17.10 | 35.40 | 67.23 | ||
| de-nl-ru | MTInstruct | overall | 8.38 | 22.75 | 59.93 | enxx | 17.05 | 32.02 | 69.26 |
| xxen | 25.13 | 45.02 | 76.29 | ||||||
| seenseen | 14.52 | 27.25 | 70.48 | enseen | 17.60 | 29.87 | 73.81 | ||
| seenunseen | 6.14 | 22.82 | 54.75 | enunseen | 16.50 | 34.17 | 64.70 | ||
| unseenseen | 7.56 | 19.22 | 61.99 | seenen | 25.73 | 47.07 | 77.52 | ||
| unseenunseen | 6.85 | 23.45 | 54.07 | unseenen | 24.53 | 42.97 | 75.06 | ||
| MT+Align | overall | 8.86 | 23.30 | 60.70 | enxx | 16.63 | 31.73 | 68.79 | |
| xxen | 25.62 | 45.37 | 76.45 | ||||||
| seenseen | 14.77 | 27.80 | 71.07 | enseen | 15.80 | 28.47 | 72.35 | ||
| seenunseen | 6.31 | 23.08 | 54.81 | enunseen | 17.47 | 35.00 | 65.24 | ||
| unseenseen | 8.61 | 20.24 | 63.81 | seenen | 25.90 | 47.13 | 77.47 | ||
| unseenunseen | 7.15 | 23.70 | 54.51 | unseenen | 25.33 | 43.60 | 75.43 | ||
| ar-de-fr-nl-ru-zh | MTInstruct | overall | 11.79 | 26.36 | 63.22 | enxx | 21.18 | 35.52 | 70.86 |
| xxen | 28.35 | 48.00 | 77.30 | ||||||
| seenseen | 22.68 | 35.32 | 76.39 | enseen | 26.20 | 37.77 | 78.22 | ||
| seenunseen | 7.10 | 24.50 | 55.18 | enunseen | 16.17 | 33.27 | 63.50 | ||
| unseenseen | 12.56 | 24.74 | 68.83 | seenen | 31.97 | 52.93 | 79.72 | ||
| unseenunseen | 6.78 | 22.62 | 53.69 | unseenen | 24.73 | 43.07 | 74.88 | ||
| MT+Align | overall | 12.13 | 26.65 | 63.23 | enxx | 21.33 | 35.65 | 70.99 | |
| xxen | 28.60 | 48.27 | 77.49 | ||||||
| seenseen | 23.67 | 36.53 | 76.89 | enseen | 26.30 | 37.63 | 78.25 | ||
| seenunseen | 7.27 | 24.32 | 54.96 | enunseen | 16.37 | 33.67 | 63.73 | ||
| unseenseen | 12.92 | 25.29 | 69.10 | seenen | 32.03 | 53.07 | 79.93 | ||
| unseenunseen | 6.68 | 22.30 | 53.19 | unseenen | 25.17 | 43.47 | 75.05 | ||
4 Evaluation and Analysis
This section reports BLEU Papineni et al. (2002); Post (2018), chrF++ Popović (2015), and COMET Rei et al. (2020)999COMET scores do not currently support Limburgish (li), Occitan (oc), Tajik (tg), Turkmen (tk), and Tatar (tt) among the 24 languages in the BLOOMZ+24 setting. Thus, we report the average COMET scores for the remaining 19 languages. scores for respective experimental configurations. We further characterized of the degree to which intermediate embeddings were language-agnostic after fine-tuning.
4.1 BLOOMZ+24 Results
Tab. 1 shows the scores for the unmodified BLOOMZ models, as well as BLOOMZ+24 under MTInstruct, AlignInstruct, and the three distinct curricula. Non-trivial improvements in all metrics were evident for BLOOMZ+24 under MTInstruct. This suggests that MTInstruct can induce translation capabilities in unseen languages. Applying AlignInstruct and MTInstruct via the curricula further showed better scores than the baselines, suggesting the role of AlignInstruct as complementing MTInstruct. AlignMT was an exception, performing similarly to MTInstruct. This may indicate the effect of AlignInstruct depends on its cadence relative to MTInstruct in a curriculum.
Superior OPUS and Flores scores under the xxen direction were evident, compared to the reverse direction, enxx. This suggests that our treatments induced understanding capabilities more than generative ones. This may be attributed to the fact that BLOOMZ had significant exposure to English, and that we used English-centric corpora. Finally, we noted the inferior performance of Flores than OPUS. This speaks to the challenge of instilling out-of-domain translation abilities in unseen languages. Our future work will focus on enhancing the domain generalization capabilities of LLM fine-tuning in MT tasks.
Moreover, we reported the language-wise scores in Tab. 2. Specifically, in the “en-xx” direction, 11 languages showed statistically significant Koehn (2004) improvements, and only 2 decreased significantly. In the “xx-en” direction, the improvements were more pronounced, with 18 languages improving significantly (most by over 1 BLEU point) and 3 decreasing significantly. The average improvement for “en-xx” was 0.74, which was substantial, especially given the limited volume of parallel data available for each language. The smaller average increase in “xx-en” can be attributed to a large decrease in one language (gd), likely due to limited training data (which can be potentially addressed with oversampling). The significantly enhanced performance in most individual languages underscores the effectiveness of our proposed methods.
4.2 Assessing AlignInstruct Variants
From Tab. 3, we observed the objectives with AlignInstruct consistently outperformed those with HintInstruct or ReviseInstruct across metrics and model sizes. Namely, easy, discriminative instructions, rather than hard, generative ones, may be preferred for experiments under similar data constraints. The low-resource constraint likely made MTInstruct more sensitive to the difficulty of its accompanying tasks.
Further, combining more than two instruction tuning tasks simultaneously did not guarantee consistent improvements, see Tab. 4. Notably, MT+Align either outperformed or matched the performance of other objective configurations. While merging multiple instruction tuning tasks occasionally resulted in superior BLEU and chrF++ scores for OPUS xxen, it fell short in COMET scores compared to MT+Align. This indicated that while such configurations might enhance word-level translation quality, as reflected by BLEU and chrF++ scores, due to increased exposure to cross-lingual word alignments, MT+Align better captured the context of the source sentence as reflected by COMET scores. Overall, these instruction tuning tasks did not demonstrate significant synergistic effects for fine-tuning for unseen languages.
4.3 BLOOMZ+3 Zero-shot Evaluation
Tab. 5 reports the results of the two settings, de-nl-ru and ar-de-fr-nl-ru-zh. Results of MT+Align+Hint+Revise and pivot-based translation are reported in App. C and H. In the de-nl-ru setting, where BLOOMZ was fine-tuned with the three unseen languages, we noticed MT+Align consistently outperformed the MTInstruct baseline across all evaluated zero-shot directions. Notably, MT+Align enhanced the translation quality for unseenseen and seenunseen directions compared to w/o fine-tuning and MTInstruct, given that the model was solely fine-tuned on de, nl, and ru data. This suggested AlignInstruct not only benefits the languages supplied in the data but also has a positive impact on other languages through cross-lingual alignment supervision. In terms of supervised directions involving English, we noticed performance improvements associated with unseen languages, and regression in seen ones. The regression may be attributed to forgetting for the absence of seen languages in fine-tuning data. Indeed, continuous exposure to English maintained the translation quality for seenen. As LoRA is modular, the regression can be mitigated by detaching the LoRA parameters for seen languages.
The ar-de-fr-nl-ru-zh setting yielded a consistently higher translation quality across all directions when compared with the de-nl-ru setting. This improvement was expected, as all the six languages were included. Translation quality improved for when generating seen languages under the zero-shot scenario. However, the same observation cannot be made for unseen languages. This phenomenon underscored the effectiveness of AlignInstruct in enhancing translation quality for BLOOMZ’s supported languages, but suggested limitations for unseen languages when mixed with supported languages in zero-shot scenarios. In the supervised directions, we found all translation directions surpassed the performance of the MTInstruct baseline. This highlighted the overall effectiveness of AlignInstruct in enhancing translation quality across a range of supervised directions.
4.4 How did MTInstruct and AlignInstruct Impact BLOOMZ’s Representations?
This section analyzed the layer-wise cosine similarities between the embeddings of parallel sentences to understand the changes in internal representations after fine-tuning. The parallel sentences were prepared from the English-centric validation datasets. We then mean-pool the outputs at each layer as sentence embeddings and compute the cosine similarities, as illustrated in Fig. 3. Results for BLOOMZ+3 are discussed in App. D.
We observed that, after MTInstruct fine-tuning, the cosine similarities rose in nearly all layers (1, Fig. 3). This may be interpreted as enhanced cross-lingual alignment, and as indicating the acquisition of translation capabilities. Upon further combination with AlignInstruct (2, Fig. 3), the degree of cross-lingual alignment rose in the early layers (layers 4 - 7) then diminished in the final layers (layers 29 & 30). This pattern aligned with the characteristics of encoder-decoder multilingual NMT models, where language-agnostic encoder representations with language-specific decoder representations improve multilingual NMT performance Liu et al. (2021); Wu et al. (2021); Mao et al. (2023). This highlights the beneficial impact of AlignInstruct.
5 Related Work
Prompting LLMs for MT LLMs have shown good performance for multilingual MT through few-shot in-context learning (ICL) Jiao et al. (2023). Agrawal et al. (2023) and Zhang et al. (2023a) explored strategies to compose better examples for ICL for XGLM-7.5B Lin et al. (2022) and GLM-130B Zeng et al. (2023). Ghazvininejad et al. (2023), Peng et al. (2023), and Moslem et al. (2023) claimed that dictionary-based hints and domain-specific style information can improve prompting OPT Zhang et al. (2022), GPT-3.5 Brown et al. (2020), and BLOOM Scao et al. (2022) for MT. He et al. (2023) used LLMs to mine useful knowledge for prompting GPT-3.5 for MT.
Fine-tuning LLMs for MT ICL-based methods do not support languages unseen during pre-training. Current approaches address this issue via fine-tuning. Zhang et al. (2023b) explored adding new languages to LLaMA Touvron et al. (2023a) with interactive translation task for unseen high-resource languages. However, similar task datasets are usually not available for most unseen, low-resource languages. Li et al. (2023) and Xu et al. (2023a) showed multilingual fine-tuning with translation instructions can improve the translation ability in supported languages. Our study extended their finding to apply in the context of unseen, low-resource languages. In parallel research, Yang et al. (2023) undertook MT instruction fine-tuning in a massively multilingual context for unseen languages. However, their emphasis was on fine-tuning curriculum based on resource availability of languages, whereas we exclusively centered on low-resource languages and instruction tuning tasks.
6 Conclusion
In this study, we introduced AlignInstruct for enhancing the fine-tuning of LLMs for MT in unseen, low-resource languages while limiting the use of additional training corpora. Our multilingual and zero-shot findings demonstrated the strength of AlignInstruct over the MTInstruct baseline and other instruction variants. Our future work pertains to exploring using large monolingual corpora of unseen languages for MT and refining the model capability to generalize across diverse MT prompts.
Limitations
Multilingual LLMs In this study, our investigations were confined to the fine-tuning of BLOOMZ models with sizes of 1.1B, 3B, and 7.1B. We did not experiment with the 175B BLOOMZ model due to computational resource constraints. However, examining this model could provide valuable insights into the efficacy of our proposed techniques. Additionally, it would be instructive to experiment with other recent open-source multilingual LLMs, such as mGPT Shliazhko et al. (2022) and LLaMa2 Touvron et al. (2023b).
PEFT Methods and Adapters As discussed in the BLOOM+1 paper Yong et al. (2023), alternative PEFT techniques, such as (IA)3 Liu et al. (2022), have the potential to enhance the adaptation performance of LLM pre-training for previously unseen languages. These approaches are worth exploring for MT fine-tuning in such languages, in addition to the LoRA methods employed in this study. Furthermore, our exploration was limited to fine-tuning multiple languages using shared additional parameters. Investigating efficient adaptation through the use of the mixture of experts (MoE) approach for MT tasks Fan et al. (2021); Costa-jussà et al. (2022); Mohammadshahi et al. (2022); Koishekenov et al. (2023); Xu et al. (2023b) presents another intriguing avenue for LLM fine-tuning.
Instruction Fine-tuning Data Another limitation of our study is that we exclusively explored MT instruction fine-tuning using fixed templates to create MT and alignment instructions. Investigating varied templates (either manually Yang et al. (2023) or automatically constructed Zhou et al. (2023)) might enhance the fine-tuned MT model’s ability to generalize across different MT task descriptions. Additionally, leveraging large monolingual corpora in unseen languages could potentially enhance the effectiveness of monolingual instructions for MT downstream tasks, offering further insights beyond the resource-constrained scenarios examined in this work. Furthermore, the creation and utilization of instruction tuning datasets, akin to xP3 Muennighoff et al. (2023), for unseen, low-resource languages could potentially amplify LLMs’ proficiency in following instructions in such languages. Zhu et al. (2023b) has investigated multilingual instruction tuning datasets. However, the scalability of such high-quality datasets to thousands of low-resource languages still remains to be addressed.
Comparison with the State-of-the-art Multilingual NMT Models In this study, we refrained from contrasting translations in low-resource languages with best-performing multilingual NMT models like NLLB-200 Costa-jussà et al. (2022), as our primary objective centered on enhancing the MTInstruct baseline through improved cross-lingual alignment within LLMs, rather than delving into the best combination of techniques for MT fine-tuning in LLMs. In future exploration, our methods can potentially be integrated with the MT fine-tuning paradigm proposed by the concurrent work of Xu et al. (2023a), paving the way for elevating the state-of-the-art translation quality using LLMs.
Appendix A Training Data Statistics
Training data statistics of BLOOMZ+24 are shown in Tab. 6. Several selected languages involved previously unseen scripts by BLOOMZ, but such fine-tuning is practical as BLOOMZ is a byte-level model with the potential to adapt to any language. Note that our proposed methods can be applied to any byte-level generative LLMs.
Appendix B Implementation Details
We employed V100 GPUs for the BLOOMZ+24 and V100 GPUs for the BLOOMZ+3 experiments. The batch sizes were configured at sentences for BLOOMZ-7b1 and sentences for both BLOOMZ-3b and BLOOMZ-1b1, per GPU device. We configured LoRA with a rank of , an alpha of , and a dropout of . Consequently, the BLOOMZ-7b1, BLOOMZ-3b, and BLOOMZ-1b1 models had M, M, and M trainable parameters, respectively, constituting approximately - % of the parameters in the original models. We conducted training for epochs, ensuring a stable convergence is achieved. To facilitate this stability, we introduced a warm-up ratio of into our training process. Maximum input and output length were set as . for HintInstruct was set as at most. Additionally, we used mixed precision training Micikevicius et al. (2018) to expedite computation using DeepSpeed Rasley et al. (2020). We tuned the optimal learning rate for each individual experiment according to validation loss. We conducted all experiments once due to computational resource constraints and reported the average scores across all languages.
Appendix C Results of MT+Align+Hint+Revise for BLOOMZ+3
We present the results in Tab. 7. Co-referencing the results in Tab. 5, compared with MT+Align, we observed a clear advantage for the MT+Align+Hint+Revise setting in supervised directions involving English (enseen and seenen) in the ar-fr-de-nl-ru-zh setting. This result suggested that AlignInstruct’s variants played a crucial role in preserving the BLOOMZ’s capabilities for supported languages. However, in all other scenarios, AlignInstruct alone proved sufficient to enhance the performance beyond the MTInstruct baseline, but hard to achieve further improvements with additional instructions.
Appendix D Representation Change of BLOOMZ+3
The representation change observed in de-nl-ru was consistent with the findings presented in Sec. 4.4, which highlighted an initial increase in cross-lingual alignment in the early layers, followed by a decrease in the final layers. When mixing fine-tuning data with supported languages, the changes exhibited more intricate patterns. As illustrated by ar-fr-zh in ar-de-fr-nl-ru-zh in Fig. 4, sentence alignment declined after MTInstruct fine-tuning but elevated after further combining with AlignInstruct. We leave the interpretation of this nuanced behavior in future work.
Appendix E Examples of HintInstruct and ReviseInstruct
We illustrated examples of HintInstruct and ReviseInstruct in Fig. 5.
| Objective | OPUS enxx | OPUS xxen | Flores enxx | Flores xxen | ||||||||
| BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | |
| MTInstruct | 11.54 | 25.33 | 64.54 | 18.59 | 33.25 | 69.84 | 3.30 | 17.10 | 40.58 | 11.37 | 27.14 | 56.33 |
| MT+Mono-full | 9.89 | 22.42 | 62.52 | 15.43 | 29.04 | 66.64 | 3.00 | 16.68 | 40.49 | 10.26 | 25.15 | 54.17 |
| MT+Mono-half | 10.23 | 22.45 | 62.22 | 15.51 | 29.65 | 67.29 | 3.18 | 16.91 | 40.57 | 10.66 | 26.15 | 54.80 |
| MT+Mono-full+Align | 10.15 | 22.35 | 62.22 | 15.72 | 29.86 | 67.70 | 3.07 | 16.59 | 40.78 | 10.61 | 25.58 | 55.17 |
| MT+Mono-half+Align | 10.09 | 22.61 | 62.98 | 16.00 | 30.34 | 67.96 | 3.10 | 16.75 | 40.70 | 10.79 | 26.27 | 55.40 |
| MT+Mono-full+Align+Hint+Revise | 10.33 | 23.04 | 63.19 | 17.16 | 31.61 | 68.26 | 3.23 | 16.70 | 40.90 | 10.98 | 26.18 | 55.50 |
| MT+Mono-half+Align+Hint+Revise | 10.62 | 23.10 | 62.92 | 17.32 | 31.80 | 68.56 | 3.20 | 16.93 | 41.00 | 11.09 | 26.77 | 55.99 |
Appendix F Assessing Monolingual Instructions
New language capabilities may be induced through continual pre-training on monolingual next-word prediction tasks Yong et al. (2023). The coherence of the generated sentences is crucial in MT Wang et al. (2020); Liu et al. (2020), especially when the target languages are unseen and low-resource. We examined the significance of this approach in fostering the translation quality. We reused the same parallel corpora to avoid introducing additional monolingual datasets.
Given a monolingual sentence, , with length in an unseen language . The LLM is incrementally trained on the following task:
Input: “Given the context, complete the following sentence: ,”
Output: “.”
We conducted experiments with two MonoInstruct settings: MonoInstruct-full, an objective to generate the entire sentence, and MonoInstruct-half for generating the latter half of the sentence given the first half, inspired by GPT Radford et al. (2018) and MASS Song et al. (2019), respectively. We reported the MonoInstruct results in Tab. 8. Firstly, we observed that fine-tuning MTInstruct in conjunction with either MonoInstruct-full or MonoInstruct-half harms the MT performance, which could be attributed to the inherent difficulty of monolingual instruction tasks and the limited amount of monolingual data. We found that the simpler MT+Mono-half yielded better results than MT+Mono-full as richer contexts were provided. However, MonoInstruct still did not improve the MTInstruct baseline. Secondly, further combining MonoInstrcut with AlignInstruct variants yielded improvements compared with MT+Mono-full (or half), but underperformed the MTInstruct baseline. This suggested that improving MT performance with monolingual instructions is challenging without access to additional monolingual data.
Appendix G Inference using Different MT Prompts
We investigated the performance of fine-tuned models when using various MT prompts during inference, aiming to understand models’ generalization capabilities with different test prompts. We examined five MT prompts for the fine-tuned models of BLOOMZ-7b1, following Zhang et al. (2023a), which are presented in Tab. 9. The results, showcased in Tab. 10, revealed that in comparison to the default prompt used during fine-tuning, the translation performance tended to decline when using other MT prompts. We observed that MT+Align consistently surpasses MTInstruct for xxen translations, though the results were mixed for enxx directions. Certain prompts, such as PROMPT-3 and PROMPT-4, exhibited a minor performance drop, while others significantly impacted translation quality. These findings underscored the need for enhancing the models’ ability to generalize across diverse MT prompts, potentially by incorporating a range of MT prompt templates during the fine-tuning process, as stated in the Limitations section.
Appendix H Zero-shot Translation using English as Pivot
| Prompt | Definition |
| PROMPT-default | Translate from to . |
| PROMPT-1 | : . |
| PROMPT-2 | . |
| PROMPT-3 | Translate to . |
| PROMPT-4 | Translate from to . |
| PROMPT-5 | Translate to . |
| Prompt | Objective | enxx | xxen | ||||
| BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | ||
| PROMPT-default | MTInstruct | 11.54 | 25.33 | 64.54 | 18.59 | 33.25 | 69.84 |
| MT+Align | 12.28 | 26.17 | 65.54 | 18.72 | 34.02 | 70.69 | |
| PROMPT-1 | MTInstruct | 5.29 | 11.31 | 50.20 | 7.87 | 20.08 | 57.46 |
| MT+Align | 5.30 | 11.38 | 50.95 | 8.93 | 20.77 | 58.38 | |
| PROMPT-2 | MTInstruct | 2.20 | 6.68 | 45.56 | 7.15 | 19.08 | 57.22 |
| MT+Align | 1.91 | 5.35 | 43.84 | 7.61 | 18.80 | 56.76 | |
| PROMPT-3 | MTInstruct | 10.59 | 22.69 | 62.65 | 15.85 | 29.93 | 67.59 |
| MT+Align | 9.20 | 20.80 | 60.96 | 16.17 | 30.58 | 68.70 | |
| PROMPT-4 | MTInstruct | 8.67 | 20.73 | 61.50 | 15.20 | 28.95 | 66.61 |
| MT+Align | 8.91 | 20.53 | 61.64 | 16.25 | 30.67 | 67.94 | |
| PROMPT-5 | MTInstruct | 6.61 | 14.55 | 55.99 | 10.88 | 22.41 | 61.40 |
| MT+Align | 6.02 | 12.28 | 52.42 | 11.83 | 23.85 | 62.09 | |
| MTInstruct | BLEU | chrF++ | COMET | MT+Align | BLEU | chrF++ | COMET |
| overall | 11.79 | 26.36 | 63.22 | overall | 12.13 | 26.65 | 63.23 |
| seenseen | 22.68 | 35.32 | 76.39 | seenseen | 23.67 | 36.53 | 76.89 |
| seenunseen | 7.10 | 24.50 | 55.18 | seenunseen | 7.27 | 24.32 | 54.96 |
| unseenseen | 12.56 | 24.74 | 68.83 | unseenseen | 12.92 | 25.29 | 69.10 |
| unseenunseen | 6.78 | 22.62 | 53.69 | unseenunseen | 6.68 | 22.30 | 53.19 |
| MTInstruct with English pivot | BLEU | chrF++ | COMET | MT+Align with English pivot | BLEU | chrF++ | COMET |
| overall | 12.99 | 28.01 | 65.38 | overall | 13.25 | 28.30 | 65.57 |
| seenseen | 23.10 | 35.30 | 76.30 | seenseen | 23.48 | 35.57 | 76.43 |
| seenunseen | 9.00 | 27.67 | 59.54 | seenunseen | 9.28 | 28.03 | 59.73 |
| unseenseen | 13.18 | 24.98 | 68.77 | unseenseen | 13.36 | 25.22 | 68.94 |
| unseenunseen | 8.57 | 25.77 | 58.17 | unseenunseen | 8.83 | 26.07 | 58.42 |
Pivot translation serves as a robust technique for zero-shot translation, especially given that we used English-centric data during fine-tuning. In Tab. 11, we present results that utilize English as an intermediary pivot for translations between non-English language pairs. Our findings indicated that employing the English pivot typically yielded an enhancement of approximately 1.1 - 1.2 BLEU points compared to direct translations in zero-shot directions when fine-tuning BLOOMZ. When contrasting the MTInstruct baseline with our proposed MT+Align, we observed that combining AlignInstruct consistently boosted performance in pivot translation scenarios.
Appendix I Per Language Result Details of BLOOMZ+24 and BLOOMZ+3
We present per language detailed results of original BLOOMZ-7b1 and fine-tuned BLOOMZ-7b1 models in Tab. 12, 13, 14, 15, 16, 17, 18, 19, respectively for the BLOOMZ+24 and BLOOMZ+3 settings.
| Language | OPUS enxx | OPUS xxen | Flores enxx | Flores xxen | ||||||||
| BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | |
| af | 3.8 | 13.2 | 56.38 | 7.6 | 22.0 | 59.14 | 2.6 | 14.9 | 33.60 | 20.1 | 38.0 | 65.61 |
| am | 0.1 | 0.3 | 33.17 | 0.5 | 8.3 | 43.57 | 0.3 | 0.6 | 30.65 | 1.9 | 12.6 | 46.24 |
| be | 4.2 | 5.1 | 47.26 | 7.3 | 17.5 | 48.57 | 0.4 | 3.3 | 31.58 | 4.2 | 22.3 | 49.27 |
| cy | 2.7 | 10.5 | 53.21 | 6.2 | 16.0 | 53.25 | 1.2 | 11.2 | 34.17 | 6.0 | 20.3 | 53.45 |
| ga | 1.2 | 10.6 | 42.85 | 4.0 | 16.4 | 46.05 | 1.2 | 11.6 | 33.94 | 5.5 | 19.6 | 46.97 |
| gd | 9.3 | 16.0 | 51.40 | 47.6 | 55.9 | 59.30 | 1.2 | 11.2 | 36.28 | 4.2 | 18.8 | 43.73 |
| gl | 4.5 | 25.6 | 64.93 | 17.2 | 36.7 | 66.07 | 13.4 | 38.5 | 74.77 | 51.0 | 67.8 | 85.77 |
| ha | 0.1 | 5.4 | 38.42 | 0.3 | 11.2 | 42.58 | 1.5 | 10.2 | 35.77 | 6.9 | 18.9 | 47.37 |
| ka | 0.3 | 1.9 | 31.97 | 0.6 | 9.2 | 44.48 | 0.4 | 1.4 | 28.81 | 2.4 | 17.0 | 47.57 |
| kk | 4.3 | 4.9 | 50.51 | 5.1 | 14.2 | 51.51 | 0.5 | 1.6 | 33.66 | 5.1 | 19.8 | 51.40 |
| km | 2.8 | 4.5 | 51.68 | 3.9 | 11.1 | 50.40 | 0.8 | 2.9 | 39.56 | 5.6 | 16.2 | 50.42 |
| ky | 10.0 | 10.6 | 54.23 | 10.3 | 24.0 | 55.99 | 0.6 | 1.6 | 30.19 | 3.8 | 17.9 | 48.05 |
| li | 6.6 | 16.2 | - | 5.9 | 24.8 | - | 2.0 | 14.9 | - | 9.8 | 29.8 | - |
| my | 1.8 | 2.4 | 45.44 | 3.0 | 5.0 | 48.33 | 0.4 | 0.8 | 29.58 | 1.0 | 3.7 | 44.15 |
| nb | 5.8 | 18.2 | 57.01 | 13.9 | 33.0 | 56.37 | 3.9 | 19.3 | 46.74 | 19.8 | 40.3 | 63.56 |
| nn | 6.3 | 18.6 | 62.33 | 8.9 | 25.3 | 56.28 | 3.7 | 19.7 | 41.75 | 16.9 | 37.5 | 62.37 |
| oc | 6.0 | 13.6 | - | 5.1 | 18.6 | - | 9.6 | 33.6 | - | 53.0 | 68.5 | - |
| si | 0.6 | 2.0 | 41.84 | 1.6 | 9.4 | 48.58 | 0.5 | 1.4 | 28.08 | 1.6 | 9.1 | 42.67 |
| tg | 0.4 | 1.4 | - | 1.1 | 11.8 | - | 0.4 | 1.5 | - | 3.3 | 18.0 | - |
| tk | 7.9 | 10.6 | - | 5.3 | 13.0 | - | 0.7 | 8.7 | - | 4.2 | 20.1 | - |
| tt | 0.0 | 1.0 | - | 0.2 | 13.3 | - | 0.3 | 1.4 | - | 4.2 | 20.2 | - |
| ug | 0.0 | 0.4 | 32.44 | 0.3 | 11.2 | 45.69 | 0.3 | 0.9 | 31.34 | 3.0 | 16.5 | 48.99 |
| uz | 0.7 | 2.1 | 35.94 | 1.0 | 12.8 | 41.86 | 1.5 | 11.5 | 40.65 | 3.1 | 18.7 | 49.43 |
| yi | 7.3 | 16.5 | 57.47 | 4.0 | 23.0 | 63.91 | 0.7 | 1.7 | 33.22 | 2.1 | 15.6 | 41.87 |
| avg. | 3.61 | 8.82 | 47.81 | 6.70 | 18.49 | 51.68 | 2.00 | 9.35 | 36.54 | 9.95 | 24.47 | 52.05 |
| Language | OPUS enxx | OPUS xxen | Flores enxx | Flores xxen | ||||||||
| BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | |
| af | 25.0 | 41.4 | 71.05 | 38.5 | 52.3 | 78.94 | 10.1 | 31.0 | 45.42 | 33.9 | 51.1 | 72.66 |
| am | 3.0 | 12.8 | 59.55 | 3.4 | 19.8 | 59.71 | 0.2 | 5.2 | 42.97 | 1.4 | 16.0 | 49.47 |
| be | 8.9 | 14.9 | 55.16 | 14.0 | 24.9 | 62.37 | 0.7 | 12.3 | 30.90 | 3.7 | 21.0 | 49.99 |
| cy | 20.2 | 38.0 | 71.55 | 33.2 | 49.3 | 77.72 | 5.0 | 20.3 | 38.38 | 13.1 | 30.2 | 57.47 |
| ga | 15.6 | 37.1 | 63.87 | 29.2 | 49.1 | 75.94 | 3.7 | 21.2 | 39.17 | 12.5 | 30.3 | 57.53 |
| gd | 13.1 | 24.7 | 62.14 | 66.0 | 69.6 | 77.70 | 2.2 | 19.6 | 40.75 | 7.1 | 22.3 | 50.05 |
| gl | 16.9 | 37.6 | 70.62 | 24.7 | 43.6 | 75.62 | 21.9 | 45.2 | 77.26 | 46.6 | 64.5 | 86.86 |
| ha | 12.3 | 32.7 | 71.75 | 10.0 | 29.8 | 64.51 | 1.9 | 17.1 | 49.24 | 6.8 | 22.1 | 48.81 |
| ka | 4.6 | 18.1 | 67.39 | 10.0 | 24.3 | 60.50 | 0.3 | 6.8 | 27.46 | 1.5 | 14.9 | 46.10 |
| kk | 12.6 | 19.5 | 66.07 | 14.6 | 28.2 | 71.80 | 0.8 | 13.0 | 35.76 | 3.9 | 19.7 | 52.24 |
| km | 19.7 | 25.2 | 63.24 | 13.9 | 32.1 | 75.02 | 0.5 | 12.3 | 35.60 | 6.2 | 22.4 | 56.45 |
| ky | 16.0 | 20.5 | 66.27 | 21.1 | 33.8 | 73.06 | 0.9 | 12.7 | 36.10 | 3.0 | 17.5 | 50.40 |
| li | 13.5 | 32.8 | - | 21.3 | 35.7 | - | 3.3 | 19.9 | - | 14.6 | 31.4 | - |
| my | 6.2 | 14.3 | 58.04 | 5.2 | 15.6 | 63.65 | 0.2 | 12.9 | 40.37 | 1.3 | 12.7 | 48.38 |
| nb | 12.7 | 30.4 | 63.27 | 22.2 | 42.1 | 76.74 | 7.9 | 28.4 | 44.15 | 25.6 | 44.3 | 72.56 |
| nn | 18.3 | 38.0 | 77.18 | 27.1 | 47.7 | 81.80 | 7.3 | 25.7 | 45.35 | 24.3 | 42.9 | 70.06 |
| oc | 10.0 | 20.0 | - | 13.4 | 27.1 | - | 8.0 | 27.5 | - | 46.9 | 63.5 | - |
| si | 5.2 | 21.4 | 68.16 | 11.5 | 26.4 | 70.79 | 0.9 | 12.9 | 41.73 | 3.7 | 19.2 | 57.41 |
| tg | 5.5 | 22.0 | - | 8.0 | 25.9 | - | 1.1 | 15.8 | - | 3.1 | 19.6 | - |
| tk | 24.4 | 26.7 | - | 30.4 | 37.8 | - | 0.7 | 10.8 | - | 3.9 | 18.8 | - |
| tt | 1.9 | 17.6 | - | 3.6 | 19.6 | - | 0.4 | 13.7 | - | 1.6 | 14.3 | - |
| ug | 1.2 | 19.7 | 49.76 | 4.2 | 21.2 | 61.34 | 0.4 | 12.9 | 35.88 | 1.7 | 16.7 | 50.29 |
| uz | 3.1 | 18.2 | 62.12 | 5.7 | 22.0 | 61.12 | 0.5 | 3.6 | 34.67 | 3.9 | 18.8 | 50.32 |
| yi | 7.1 | 24.3 | 59.13 | 14.9 | 20.2 | 58.66 | 0.3 | 9.5 | 29.77 | 2.5 | 17.2 | 43.27 |
| avg. | 11.54 | 25.33 | 64.54 | 18.6 | 33.25 | 68.84 | 3.30 | 17.10 | 40.58 | 11.37 | 27.14 | 56.33 |
| Language | OPUS enxx | OPUS xxen | Flores enxx | Flores xxen | ||||||||
| BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | BLEU | chrF++ | COMET | |
| af | 25.0 | 41.9 | 70.72 | 36.9 | 52.2 | 78.68 | 10.6 | 31.9 | 45.84 | 33.5 | 51.1 | 72.84 |
| am | 3.4 | 13.2 | 60.62 | 4.9 | 22.8 | 62.43 | 0.3 | 5.4 | 44.20 | 1.4 | 16.4 | 51.05 |
| be | 8.3 | 14.5 | 55.23 | 13.9 | 25.1 | 62.72 | 0.8 | 12.5 | 30.93 | 3.6 | 20.6 | 49.14 |
| cy | 20.6 | 39.0 | 71.73 | 33.8 | 49.4 | 77.55 | 4.7 | 20.3 | 38.70 | 14.6 | 31.5 | 58.34 |
| ga | 17.6 | 39.3 | 65.76 | 32.6 | 52.7 | 77.49 | 3.4 | 21.4 | 39.99 | 13.6 | 31.6 | 58.73 |
| gd | 15.6 | 27.2 | 62.09 | 48.1 | 55.4 | 75.90 | 2.3 | 20.3 | 40.81 | 7.4 | 22.0 | 49.99 |
| gl | 17.1 | 37.2 | 70.85 | 24.4 | 43.3 | 75.90 | 21.7 | 44.9 | 77.09 | 45.6 | 63.5 | 86.60 |
| ha | 14.6 | 35.0 | 73.34 | 11.4 | 31.3 | 65.69 | 1.9 | 17.3 | 50.88 | 7.4 | 22.5 | 49.57 |
| ka | 4.9 | 18.9 | 67.54 | 10.5 | 25.3 | 61.27 | 0.3 | 6.9 | 27.61 | 2.1 | 16.0 | 47.04 |
| kk | 12.3 | 19.3 | 65.73 | 15.6 | 28.0 | 71.01 | 0.9 | 13.0 | 35.86 | 4.1 | 19.8 | 52.43 |
| km | 20.4 | 26.5 | 63.38 | 14.4 | 35.2 | 75.62 | 0.6 | 12.5 | 35.44 | 7.1 | 22.9 | 57.81 |
| ky | 15.8 | 19.6 | 64.74 | 23.3 | 35.8 | 74.70 | 0.9 | 13.3 | 36.71 | 2.9 | 17.4 | 50.06 |
| li | 13.2 | 29.4 | - | 22.3 | 38.2 | - | 3.1 | 19.7 | - | 12.5 | 28.7 | - |
| my | 7.6 | 15.4 | 58.84 | 6.3 | 18.0 | 66.45 | 0.3 | 13.3 | 40.97 | 1.2 | 14.4 | 50.79 |
| nb | 13.5 | 31.4 | 64.08 | 24.2 | 44.2 | 77.58 | 7.9 | 28.7 | 44.12 | 25.5 | 44.9 | 72.72 |
| nn | 19.0 | 38.0 | 77.61 | 28.5 | 47.7 | 81.68 | 7.0 | 26.7 | 46.14 | 25.8 | 44.1 | 70.55 |
| oc | 9.1 | 19.3 | - | 13.5 | 27.5 | - | 7.5 | 25.9 | - | 47.3 | 63.8 | - |
| si | 5.1 | 22.1 | 69.60 | 13.9 | 29.1 | 72.51 | 1.1 | 13.1 | 43.01 | 5.6 | 22.7 | 61.89 |
| tg | 6.6 | 23.7 | - | 8.8 | 27.2 | - | 0.9 | 15.6 | - | 3.4 | 19.9 | - |
| tk | 27.2 | 26.2 | - | 31.2 | 38.7 | - | 0.7 | 11.4 | - | 3.8 | 18.2 | - |
| tt | 2.1 | 18.6 | - | 5.0 | 21.5 | - | 0.4 | 13.3 | - | 1.5 | 13.7 | - |
| ug | 1.1 | 20.7 | 51.12 | 5.5 | 23.4 | 63.42 | 0.4 | 13.8 | 37.51 | 2.1 | 16.3 | 50.45 |
| uz | 3.5 | 18.6 | 62.09 | 7.4 | 23.3 | 62.01 | 0.2 | 1.9 | 34.50 | 3.7 | 18.2 | 50.09 |
| yi | 11.1 | 33.1 | 70.13 | 12.8 | 21.2 | 60.47 | 0.4 | 9.8 | 30.08 | 2.6 | 17.0 | 42.57 |
| avg. | 12.28 | 26.17 | 65.54 | 18.72 | 34.02 | 70.69 | 3.26 | 17.20 | 41.07 | 11.60 | 27.38 | 56.98 |
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET | |
| ar-de | 1.4 | 14.8 | 56.19 | en-ar | 11.1 | 32.4 | 75.66 | |
| ar-fr | 21.9 | 46.1 | 74.19 | en-de | 12.2 | 29.2 | 59.16 | |
| ar-nl | 0.6 | 11.2 | 56.59 | en-fr | 26.8 | 49.2 | 77.42 | |
| ar-ru | 3.1 | 6.2 | 48.41 | en-nl | 2.0 | 16.0 | 46.52 | |
| ar-zh | 18.4 | 14.4 | 73.65 | en-ru | 5.7 | 16.1 | 49.00 | |
| de-ar | 2.0 | 17.8 | 64.91 | en-zh | 22.5 | 17.0 | 77.90 | |
| de-fr | 12.0 | 33.4 | 63.45 | avg. | 13.38 | 26.65 | 64.28 | |
| de-nl | 3.7 | 17.9 | 47.30 | |||||
| de-ru | 1.3 | 11.8 | 45.53 | |||||
| de-zh | 8.9 | 7.6 | 61.52 | |||||
| fr-ar | 11.2 | 33.4 | 74.20 | BLEU | chrF++ | COMET | ||
| fr-de | 4.6 | 23.4 | 48.83 | ar-en | 26.7 | 48.4 | 78.12 | |
| fr-nl | 2.8 | 17.2 | 52.14 | de-en | 21.1 | 38.5 | 71.99 | |
| fr-ru | 3.1 | 10.4 | 45.12 | fr-en | 27.7 | 49.8 | 79.46 | |
| fr-zh | 20.9 | 17.0 | 76.20 | nl-en | 12.3 | 31.1 | 61.29 | |
| nl-ar | 1.3 | 13.2 | 59.46 | ru-en | 17.9 | 36.6 | 68.40 | |
| nl-de | 5.9 | 22.8 | 46.49 | zh-en | 24.5 | 47.9 | 77.08 | |
| nl-fr | 9.6 | 29.6 | 58.30 | avg. | 21.70 | 42.05 | 72.72 | |
| nl-ru | 0.8 | 9.0 | 42.83 | |||||
| nl-zh | 3.3 | 3.7 | 53.96 | |||||
| ru-ar | 6.5 | 25.3 | 68.38 | |||||
| ru-de | 2.0 | 17.0 | 48.06 | |||||
| ru-fr | 15.7 | 38.7 | 67.54 | |||||
| ru-nl | 0.5 | 10.5 | 46.14 | |||||
| ru-zh | 10.7 | 11.3 | 67.18 | |||||
| zh-ar | 8.6 | 29.7 | 73.47 | |||||
| zh-de | 1.6 | 17.6 | 49.90 | |||||
| zh-fr | 20.7 | 44.1 | 75.79 | |||||
| zh-nl | 0.6 | 10.4 | 48.53 | |||||
| zh-ru | 2.9 | 8.6 | 44.13 | |||||
| avg. | 6.89 | 19.14 | 57.95 | |||||
| seenseen | 16.95 | 30.78 | 74.58 | enseen | 20.13 | 32.87 | 76.99 | |
| seenunseen | 2.30 | 13.31 | 49.98 | enunseen | 6.63 | 20.43 | 51.56 | |
| unseenseen | 7.78 | 20.07 | 62.74 | seenen | 26.30 | 48.70 | 78.22 | |
| unseenunseen | 2.37 | 14.83 | 46.06 | unseenen | 17.10 | 35.40 | 67.23 |
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET | |
| ar-de | 4.7 | 20.9 | 56.43 | en-ar | 9.1 | 27.2 | 71.47 | |
| ar-fr | 20.8 | 42.5 | 71.47 | en-de | 19.8 | 36.1 | 66.53 | |
| ar-nl | 7.2 | 22.9 | 58.29 | en-fr | 23.0 | 44.5 | 74.98 | |
| ar-ru | 5.0 | 21.0 | 54.73 | en-nl | 15.5 | 36.1 | 64.76 | |
| ar-zh | 14.0 | 12.4 | 67.94 | en-ru | 14.2 | 30.3 | 62.82 | |
| de-ar | 2.4 | 16.2 | 64.53 | en-zh | 20.7 | 17.9 | 74.97 | |
| de-fr | 11.9 | 31.2 | 64.44 | avg. | 17.05 | 32.02 | 69.26 | |
| de-nl | 9.4 | 28.1 | 54.22 | |||||
| de-ru | 5.1 | 19.6 | 55.41 | |||||
| de-zh | 4.2 | 5.8 | 55.26 | |||||
| fr-ar | 10.1 | 29.1 | 70.72 | BLEU | chrF++ | COMET | ||
| fr-de | 8.6 | 27.7 | 53.77 | ar-en | 26.5 | 46.9 | 76.92 | |
| fr-nl | 10.3 | 30.1 | 57.55 | de-en | 27.0 | 44.0 | 76.97 | |
| fr-ru | 7.9 | 26.0 | 56.82 | fr-en | 27.5 | 49.0 | 78.80 | |
| fr-zh | 18.1 | 18.5 | 72.24 | nl-en | 21.8 | 41.3 | 73.99 | |
| nl-ar | 2.0 | 15.1 | 63.73 | ru-en | 24.8 | 43.6 | 74.23 | |
| nl-de | 9.7 | 28.1 | 52.58 | zh-en | 23.2 | 45.3 | 76.83 | |
| nl-fr | 13.2 | 32.3 | 65.17 | avg. | 25.13 | 45.02 | 76.29 | |
| nl-ru | 5.1 | 18.6 | 55.13 | |||||
| nl-zh | 3.0 | 5.4 | 54.34 | |||||
| ru-ar | 5.9 | 15.0 | 60.36 | |||||
| ru-de | 5.6 | 23.8 | 52.66 | |||||
| ru-fr | 17.9 | 38.4 | 68.66 | |||||
| ru-nl | 6.2 | 22.5 | 54.41 | |||||
| ru-zh | 7.5 | 13.6 | 61.40 | |||||
| zh-ar | 6.7 | 22.1 | 67.48 | |||||
| zh-de | 3.3 | 19.6 | 51.75 | |||||
| zh-fr | 17.4 | 38.9 | 73.00 | |||||
| zh-nl | 4.8 | 19.3 | 54.41 | |||||
| zh-ru | 3.5 | 17.9 | 49.02 | |||||
| avg. | 8.38 | 22.75 | 59.93 | |||||
| seenseen | 14.52 | 27.25 | 70.48 | enseen | 17.60 | 29.87 | 73.81 | |
| seenunseen | 6.14 | 22.82 | 54.75 | enunseen | 16.50 | 34.17 | 64.70 | |
| unseenseen | 7.56 | 19.22 | 61.99 | seenen | 25.73 | 47.07 | 77.52 | |
| unseenunseen | 6.85 | 23.45 | 54.07 | unseenen | 24.53 | 42.97 | 75.06 |
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET | |
| ar-de | 5.1 | 20.8 | 55.25 | en-ar | 8.4 | 26.0 | 70.45 | |
| ar-fr | 20.3 | 42.5 | 71.78 | en-de | 21.1 | 36.7 | 67.15 | |
| ar-nl | 6.4 | 21.6 | 57.48 | en-fr | 22.9 | 44.4 | 74.67 | |
| ar-ru | 5.2 | 21.5 | 55.51 | en-nl | 16.1 | 36.8 | 65.26 | |
| ar-zh | 16.0 | 14.1 | 69.55 | en-ru | 15.2 | 31.5 | 63.30 | |
| de-ar | 2.4 | 16.3 | 64.01 | en-zh | 16.1 | 15.0 | 71.93 | |
| de-fr | 13.5 | 34.3 | 66.25 | avg. | 16.63 | 31.73 | 68.79 | |
| de-nl | 9.7 | 28.0 | 55.00 | |||||
| de-ru | 5.3 | 19.6 | 55.61 | |||||
| de-zh | 7.2 | 7.3 | 60.64 | |||||
| fr-ar | 10.0 | 28.2 | 69.86 | BLEU | chrF++ | COMET | ||
| fr-de | 9.2 | 27.8 | 54.03 | ar-en | 27.1 | 47.0 | 76.54 | |
| fr-nl | 10.8 | 31.0 | 58.50 | de-en | 27.8 | 44.4 | 77.57 | |
| fr-ru | 8.6 | 26.7 | 57.07 | fr-en | 27.1 | 48.7 | 78.82 | |
| fr-zh | 15.9 | 15.8 | 70.78 | nl-en | 22.6 | 42.2 | 74.25 | |
| nl-ar | 2.2 | 15.4 | 63.47 | ru-en | 25.6 | 44.2 | 74.46 | |
| nl-de | 10.2 | 28.5 | 53.65 | zh-en | 23.5 | 45.7 | 77.04 | |
| nl-fr | 14.4 | 34.4 | 66.55 | avg. | 25.62 | 45.37 | 76.45 | |
| nl-ru | 5.3 | 19.3 | 55.53 | |||||
| nl-zh | 5.5 | 6.2 | 58.77 | |||||
| ru-ar | 6.5 | 16.0 | 62.69 | |||||
| ru-de | 6.1 | 24.3 | 52.89 | |||||
| ru-fr | 18.2 | 39.0 | 69.95 | |||||
| ru-nl | 6.3 | 22.5 | 54.36 | |||||
| ru-zh | 7.6 | 13.3 | 61.94 | |||||
| zh-ar | 8.7 | 26.5 | 70.88 | |||||
| zh-de | 3.0 | 19.5 | 50.82 | |||||
| zh-fr | 17.7 | 39.7 | 73.56 | |||||
| zh-nl | 4.4 | 19.3 | 54.20 | |||||
| zh-ru | 4.1 | 19.5 | 50.47 | |||||
| avg. | 8.86 | 23.30 | 60.70 | |||||
| seenseen | 14.77 | 27.80 | 71.07 | enseen | 15.80 | 28.47 | 72.35 | |
| seenunseen | 6.31 | 23.08 | 54.81 | enunseen | 17.47 | 35.00 | 65.24 | |
| unseenseen | 8.61 | 20.24 | 63.81 | seenen | 25.90 | 47.13 | 77.47 | |
| unseenunseen | 7.15 | 23.70 | 54.51 | unseenen | 25.33 | 43.60 | 75.43 |
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET | |
| ar-de | 6.9 | 24.7 | 58.10 | en-ar | 14.6 | 35.6 | 76.70 | |
| ar-fr | 26.2 | 48.2 | 74.96 | en-de | 20.4 | 36.0 | 65.96 | |
| ar-nl | 8.8 | 24.7 | 59.53 | en-fr | 27.9 | 50.0 | 77.65 | |
| ar-ru | 6.5 | 22.7 | 55.33 | en-nl | 14.8 | 34.8 | 63.11 | |
| ar-zh | 28.6 | 22.3 | 77.64 | en-ru | 13.3 | 29.0 | 61.43 | |
| de-ar | 3.3 | 19.8 | 68.27 | en-zh | 36.1 | 27.7 | 80.31 | |
| de-fr | 15.2 | 35.8 | 67.05 | avg. | 21.18 | 35.52 | 70.86 | |
| de-nl | 8.2 | 26.0 | 53.35 | |||||
| de-ru | 4.4 | 17.9 | 54.79 | |||||
| de-zh | 12.0 | 9.9 | 65.20 | |||||
| fr-ar | 14.2 | 35.2 | 74.84 | BLEU | chrF++ | COMET | ||
| fr-de | 8.9 | 28.4 | 53.81 | ar-en | 33.7 | 53.5 | 79.81 | |
| fr-nl | 10.1 | 29.9 | 56.92 | de-en | 27.1 | 43.9 | 77.04 | |
| fr-ru | 8.1 | 26.0 | 55.96 | fr-en | 29.6 | 51.0 | 79.60 | |
| fr-zh | 30.2 | 25.6 | 79.43 | nl-en | 22.0 | 41.4 | 73.54 | |
| nl-ar | 3.1 | 18.2 | 67.72 | ru-en | 25.1 | 43.9 | 74.05 | |
| nl-de | 10.4 | 27.7 | 52.67 | zh-en | 32.6 | 54.3 | 79.75 | |
| nl-fr | 16.9 | 37.3 | 68.46 | avg. | 28.35 | 48.00 | 77.30 | |
| nl-ru | 4.8 | 17.8 | 54.71 | |||||
| nl-zh | 8.1 | 7.0 | 63.96 | |||||
| ru-ar | 11.9 | 31.5 | 72.45 | |||||
| ru-de | 6.1 | 23.7 | 52.74 | |||||
| ru-fr | 21.2 | 42.5 | 71.71 | |||||
| ru-nl | 6.8 | 22.6 | 53.91 | |||||
| ru-zh | 21.3 | 20.7 | 74.63 | |||||
| zh-ar | 13.1 | 34.1 | 74.92 | |||||
| zh-de | 4.1 | 22.3 | 52.13 | |||||
| zh-fr | 23.8 | 46.5 | 76.54 | |||||
| zh-nl | 4.8 | 19.9 | 54.26 | |||||
| zh-ru | 5.7 | 21.9 | 50.60 | |||||
| avg. | 11.79 | 26.36 | 63.22 | |||||
| seenseen | 22.68 | 35.32 | 76.39 | enseen | 26.20 | 37.77 | 78.22 | |
| seenunseen | 7.10 | 24.50 | 55.18 | enunseen | 16.17 | 33.27 | 63.50 | |
| unseenseen | 12.56 | 24.74 | 68.83 | seenen | 31.97 | 52.93 | 79.72 | |
| unseenunseen | 6.78 | 22.62 | 53.69 | unseenen | 24.73 | 43.07 | 74.88 |
| Zero-shot | BLEU | chrF++ | COMET | Supervised | BLEU | chrF++ | COMET | |
| ar-de | 6.7 | 24.2 | 57.45 | en-ar | 15.1 | 35.8 | 76.76 | |
| ar-fr | 27.5 | 49.2 | 75.21 | en-de | 20.6 | 35.9 | 65.88 | |
| ar-nl | 8.7 | 24.8 | 59.14 | en-fr | 27.5 | 49.4 | 77.46 | |
| ar-ru | 6.7 | 21.6 | 55.04 | en-nl | 15.0 | 35.6 | 63.70 | |
| ar-zh | 30.1 | 24.4 | 78.54 | en-ru | 13.5 | 29.5 | 61.62 | |
| de-ar | 3.5 | 19.7 | 68.39 | en-zh | 36.3 | 27.7 | 80.52 | |
| de-fr | 15.4 | 35.8 | 67.81 | avg. | 21.33 | 35.65 | 70.99 | |
| de-nl | 9.6 | 27.3 | 53.74 | |||||
| de-ru | 4.7 | 17.9 | 54.23 | |||||
| de-zh | 12.0 | 9.9 | 65.40 | |||||
| fr-ar | 14.9 | 36.3 | 74.98 | BLEU | chrF++ | COMET | ||
| fr-de | 9.2 | 28.3 | 52.96 | ar-en | 33.9 | 53.7 | 79.74 | |
| fr-nl | 11.3 | 31.1 | 57.62 | de-en | 27.1 | 43.6 | 77.13 | |
| fr-ru | 8.8 | 26.2 | 56.31 | fr-en | 29.7 | 51.0 | 80.03 | |
| fr-zh | 31.1 | 26.9 | 79.93 | nl-en | 22.6 | 42.3 | 73.94 | |
| nl-ar | 3.3 | 18.5 | 68.02 | ru-en | 25.8 | 44.5 | 74.07 | |
| nl-de | 9.4 | 26.5 | 52.33 | zh-en | 32.5 | 54.5 | 80.01 | |
| nl-fr | 17.2 | 37.3 | 68.38 | avg. | 28.60 | 48.27 | 77.49 | |
| nl-ru | 4.4 | 17.1 | 53.63 | |||||
| nl-zh | 8.3 | 7.0 | 64.08 | |||||
| ru-ar | 12.4 | 32.1 | 72.40 | |||||
| ru-de | 5.7 | 22.9 | 51.90 | |||||
| ru-fr | 21.5 | 42.7 | 72.08 | |||||
| ru-nl | 6.3 | 22.1 | 53.32 | |||||
| ru-zh | 22.7 | 24.6 | 75.36 | |||||
| zh-ar | 13.9 | 35.4 | 75.68 | |||||
| zh-de | 3.6 | 21.3 | 51.32 | |||||
| zh-fr | 24.5 | 47.0 | 76.98 | |||||
| zh-nl | 4.9 | 20.3 | 54.30 | |||||
| zh-ru | 5.5 | 21.1 | 50.49 | |||||
| avg. | 12.13 | 26.65 | 63.23 | |||||
| seenseen | 23.67 | 36.53 | 76.89 | enseen | 26.30 | 37.63 | 78.25 | |
| seenunseen | 7.27 | 24.32 | 54.96 | enunseen | 16.37 | 33.67 | 63.73 | |
| unseenseen | 12.92 | 25.29 | 69.10 | seenen | 32.03 | 53.07 | 79.93 | |
| unseenunseen | 6.68 | 22.30 | 53.19 | unseenen | 25.17 | 43.47 | 75.05 |